itsmehjay:

Two numbers have a difference of 7 and a sum of -11. What are the values of x and y as an ordered pair?

itsmehjay:

@SnOwFlAkE0531

MrsHero:

@az

carmelle:

@sUpiE

itsmehjay:

@hOoDmEmEs

itsmehjay:

@SxDsOuLs e.e

itsmehjay:

why not? ;c

sxdsouls:

?

itsmehjay:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sxdsouls ? \(\color{#0cbb34}{\text{End of Quote}}\) h a l p m e h

sxdsouls:

idk this-

sxdsouls:

@supie ??

itsmehjay:

@bOrEdFr ;c

itsmehjay:

@sUpIe plz ;c

AZ:

How would you write it as an equation? The sum of two numbers is -11 The difference is 7 What would your two equations be?

itsmehjay:

idk- my mind is going blank today.......

AZ:

Call one number 'x' and the other 'y' what does sum mean? Is that adding two numbers or subtracting or multiplying or dividing?

itsmehjay:

it's adding ;-;

AZ:

Exactly, so the sum of the two numbers is -11 so x + y = -11 right?

itsmehjay:

yes yes umm -12 + 1 = -11 but it doesn't have a difference of 7

itsmehjay:

i'm so confused-

AZ:

yes, so we need to come up with the second equation so that way we can solve it what does difference mean? subtraction? division? multiplication?

itsmehjay:

subtract

AZ:

exactly, so your second equation is x - y = 7 so we have two equations x + y = -11 x - y = 7 Do you know how to solve a system of equations?

itsmehjay:

ye.. i think 0-0

AZ:

What do you get for 'x' and 'y' when you solve it?

itsmehjay:

OooHh i got ittt

itsmehjay:

so for x i got -2? and y = -9??

AZ:

That is correct